# Large Row-Constrained Supersaturated Designs for High-throughput Screening

### v1

@@ Abstract
High-throughput screening, in which multiwell plates are used to test large numbers of compounds against specific targets, is widely used across many areas of the biological sciences and most prominently in drug discovery. We propose a statistically principled approach to these screening experiments, using the machinery of supersaturated designs and the Lasso. To accommodate limitations on the number of biological entities that can be applied to a single microplate well, we present a new class of row-constrained supersaturated designs. We develop a computational procedure to construct these designs, provide some initial lower bounds on the average squared off-diagonal values of their main-effects information matrix, and study the impact of the constraint on design quality. We also show via simulation that the proposed constrained row screening method is statistically superior to existing methods and demonstrate the use of the new methodology on a real drug-discovery system.
Keywords: experimental design; biological screening; regularization
@@ 1 Introduction
High-throughput screening (HTS) of chemical or biological compounds is used extensively across a range of critical scientific enterprises. This type of screening is common in basic functional genomics studies and synthetic biology applications (Zhou et al., 2014; Zeng et al., 2020; Sarnaik et al., 2020), but arguably its highest impact application is drug discovery (Perola, 2010; Blay et al., 2020; Soheilmoghaddam et al., 2021; Aldewachi et al., 2021). The literature suggests that nearly one-third of clinical drug candidates emerge from high-throughput screening campaigns (Brown and Boström, 2018). These screens were a centerpiece in the National Cancer Institute’s Cancer Moonshot program (Wilson et al., 2020). But as liquid handling, miniaturization, and other screening technologies have improved, the number of compounds to explore has increased as well (Elkin et al., 2015; Volochnyuk et al., 2019). This, in turn, highlights the need to improve the experimental and statistical aspects of these screens. In the current work, we present a scalable statistical methodology, including a novel experimental design approach, to make high-throughput screening more flexible, cost-effective, and statistically efficient for detecting promising compounds, compared to existing methods. We do this by proposing a new type of supersaturated design which assigns a number of compounds to each well. This design is then analyzed using regularized regression. These new designs are characterized by a constraint which governs the number of compounds which can be included in each well. Using the unconditional $E(s^{2})$ supersaturated design criterion, we present a computationally efficient heuristic to construct these designs, and study them by developing lower bounds on the criterion and investigating how design quality changes as a function of the constraint. We also give evidence of the effectiveness of our approach via simulation. Finally, we demonstrate the methodology in real experimental screens used to identify an inhibitor of a metallo-$\beta$-lactamase. The prevalence of such enzymes is of great concern because they exist in bacteria that cause nosocomial infections, including in secondary bacterial infections contributing to an estimated 50% of COVID-19 deaths (Mojica et al., 2022; Chedid et al., 2021). Searching for compounds that inhibit these enzymes’ activity is an important part of the arsenal for addressing antibiotic resistance (CDC, 2019).
HTS utilizes plates with varying well capacities (e.g., 96, 384, or 1536). Each well contains cells or enzymes and is subjected to application-specific conditions. Most commonly, a single chemical compound is assigned to a well, a biological or chemical process then occurs, and a measurement is taken. By using many plates, huge compound libraries can be studied. Compounds that produce meaningful measurements are then validated by more focused experimentation. The process is called screening because the goal is to detect, among thousands of candidates, the few compounds that exhibit the most promise. Though experimental design and statistical analysis have been promoted in HTS, and sophisticated methods of normalization and bias-correction have been developed (Altman, 2005; Birmingham et al., 2009; Malo et al., 2010; Dragiev et al., 2011; Zhang, 2011; Caraus et al., 2015), the basic experimental design approach most often used is simplistic: apply each treatment to one well (Malo et al., 2006; Kainkaryam and Woolf, 2009) (or, as recommended in the literature, in triplicate; see Malo et al., 2010) and then use a z-score to determine if the result is a hit. As shown in Section 4, this approach’s advantage of simplicity comes at the cost of reduced ability to detect hits. We note the recent work using machine/deep learning to preselect the “best” compounds (e.g., Stokes et al., 2020; Cheng et al., 2024), and these approaches have become popular for drug discovery (Mayr et al., 2018; Camacho et al., 2018). Our approach is complementary to these efforts, in the sense that it might be used to create training sets for HTS machine/deep learning models.
We are not the first to propose pooling multiple compounds in each well; in fact, versions of pooling have been used successfully (see Kainkaryam and Woolf, 2009). The most aggressive version of the existing strategies is called non-adaptive pooling (Thierry-Mieg, 2006; Thierry-Mieg and Bailly, 2008; Wilson et al., 2020; Kainkaryam and Woolf, 2009), which purportedly can screen, with a high degree of accuracy, large numbers of compounds with relatively few pools. The most relevant approach uses so-called Shifted Transversal Designs (STDs) within a larger procedure called poolHiTS. This method requires that the response for each well can be identified as either a “hit” or “inert”, and also that the user specifies the expected number of hits and the expected error rate. STDs construct pools to have as little overlap between them as possible, and they use information about which pools trigger hits to back out which compounds are likely hits and which are inert. The chief advantage of STDs is that they are fast to construct, but they have several important drawbacks described in more detail in Section 4.1.2. Most basically, poolHiTS treats neither the design nor the analysis in a statistical way. Our approach, on the other hand, frames the problem as a statistically designed experiment, which generates data posited to arise from a statistical model. This allows us to use statistical estimation procedures and thereby control false positive rates with high true positive rates. Overall, we will show that our procedure is more flexible with better statistical performance than poolHiTS.
Our approach uses a supersaturated design (SSD) to guide the placement of multiple compounds in each well, where $k$ compounds are investigated using $n$ wells (with $n\leq k$) while controlling false positive rates and increasing true positive rates. SSDs—experiments for which there are more factors than runs—were first described over sixty years ago (Satterthwaite, 1959; Booth and Cox, 1962). Motivated by the desire to experiment on complex high-cost systems, SSDs have seen extensive development over the last thirty years (e.g., Lin, 1993; Wu, 1993; Nguyen, 1996; Cheng, 1997; Cheng and Tang, 2001; Butler et al., 2001; Bulutoglu and Cheng, 2004; Georgiou, 2014; Jones and Majumdar, 2014; Weese et al., 2021). In our setting, an SSD is a two-level design composed of $k$ compounds (factors; absence with level $-1$; presence with level $+1$), and $n$ wells (runs) with $k\geq n$. The SSD provides a set of instructions regarding which compounds are included in which well. The design is then analyzed assuming the main-effects statistical model
@@
where $X$ is the $n\times k$ SSD design matrix with elements $x_{ij}=\pm 1$, $\bm{\beta}=(\beta_{1},\dots,\beta_{k})^{T}$ and $\bm{\epsilon}\sim N(\mathbf{0},\sigma^{2}I)$. Because $k\geq n$, ordinary least squares cannot be used to estimate factor effects. Although forward selection approaches (Westfall et al., 1998; Abraham et al., 1999; Kelly III and Voelkel, 2000) were initially considered, regularization methods are now preferred for the analysis of SSDs. The Gauss-Dantzig selector (Candes and Tao, 2007) has been shown to be effective (Phoa et al., 2009; Marley and Woods, 2010; Weese et al., 2015), but the Lasso (Tibshirani, 1996) is closely related to the Dantzig selector (Meinshausen et al., 2007; Lounici, 2008; James et al., 2009; Asif and Romberg, 2010) and achieves similar performance (see Draguljić et al., 2014). Additionally, Stallrich et al. (2023) has recently developed an optimal screening design framework based on the Lasso. Thus, in our work we use the Lasso, which selects estimates according to
@@
A critical aspect of the proposed methodology is constructing SSDs to use as maps guiding the assignment of compounds to wells. SSDs are classically constructed using the $E(s^{2})$ criterion (Booth and Cox, 1962; Nguyen, 1996; Cheng, 1997; Bulutoglu and Cheng, 2004) to create designs with columns balanced between $-1$’s and $+1$’s, as close to mutually orthogonal as possible. We follow work that relaxes the balance requirement (Marley and Woods, 2010; Jones and Majumdar, 2014; Weese et al., 2015), leading to the unconditional $E(s^{2})$ criterion. Mathematically, let $L=[\mathbf{1},X]$, and $S=L^{{}^{\prime}}L$ with elements $s_{ij}$ and $O_{S}=\sum_{i<j}s_{ij}^{2}$. Then the unconditional $E(s^{2})$ criterion is
$UE(s^{2})=\frac{O_{S}}{k(k+1)},$ (3)
the average of the squared off-diagonal elements of $S$. In many cases $UE(s^{2})$-optimal designs can be constructed directly from known Hadamard matrices (Jones and Majumdar, 2014), but in our work we construct the designs algorithmically; see Section 2.1.
Our approach, which we call constrained row screening, or CRowS, breaks new ground in several important ways. First, it provides a statistically principled and substantially improved solution to an important problem in biology. We will demonstrate empirically in Section 4 that our approach has much better statistical properties, in terms of true and false positive rates, than existing methods. Secondly, the constraints inherent to the HTS application necessitate innovation in design construction. Specifically, though there is literature suggesting that multiple compounds can be applied to individual wells (e.g., Elkin et al., 2015; Kainkaryam and Woolf, 2008, 2009), it is also clear that in drug screening settings, and by extension high-throughput screening more broadly, there are limitations regarding the number of compounds that can be added to each well. These may be due to the size of the well, the minimum concentration required in order for a compound to be detectable, or the maximum ionic strength of the solution beyond which the compounds may precipitate. We call these limitations the row constraints, denoted by $c$, because it limits the number of $+1$’s in each row of the design. Thus, we have developed a new coordinate-exchange algorithm (Meyer and Nachtsheim, 1995) that aims to minimize $UE(s^{2})$ while respecting row constraints, and we study these new CRowS designs both theoretically and empirically as a function of $c$. Third, our work provides a significant, broad-ranging application for SSDs, which has been lacking in the literature. Despite their considerable promise (Weese et al., 2021), we only know of a handful of scattered articles that report on their use (Carpinteiro et al., 2004; Dejaegher and Vander Heyden, 2008; Jridi et al., 2015; Zarkadas and Besseris, 2023). Our work changes that, by applying these designs to a large class of biological/chemical screens and demonstrating their usefulness in real experiments. Finally, SSD research has focused largely on designs with at most dozens of factors; in contrast, CRowS designs will necessarily be larger in order to apply them to HTS. The designs we investigate in this paper have more than one hundred factors, though the proof-of-concept real experiments we report on in Section 5 are smaller. Since HTS studies routinely run into the tens of thousands and even much larger (Elkin et al., 2015), the size of the SSDs applied to this area will be limited not by the number of compounds studied, but by, in large part, logistical constraints related to the liquid handling automation (see Section 5 for a discussion) as well as computational limitations in the construction of large designs. We note that Eskridge et al. (2004) study large SSDs, and even given a nod to HTS as an application, but they don’t provide designs for $n>24$ nor do they consider row constraints or provide a real example.
The remainder of the paper is structured as follows. In Section 2, we describe row-constrained supersaturated designs in more detail, along with the algorithm we use to construct them. Section 3 provides some initial lower bounds on the $UE(s^{2})$ values for these CRowS designs, as well as an empirical study that sheds light on their behavior as a function of the severity of the row constraint. In Section 4 we use a set of extensive simulations to show the effectiveness of our proposed procedure compared to alternative methods. In Section 5 we provide more detail on the metallo-$\beta$-lactamase system, and describe the results of several screens as a real-world proof-of-concept of CRowS. We provide a discussion and conclusion in Section 6.
@@ 2 Row-Constrained Supersaturated Designs
In this section we describe the new class of supersaturated designs that have emerged from this work, along with the heuristic optimization procedure we have developed to construct them. As explained in Section 1, $UE(s^{2})$-optimal designs minimize the average of the squared off-diagonal elements of information matrix $S$. In the present setting, we propose a generalization of $UE(s^{2})$-optimal designs that impose a constraint on the number of $+1$’s in each row. Thus, using notation from Section 1, we say that an $n\times k$ design $X$ is row-constrained $UE(s^{2})$-optimal if $n\leq k$, $[X]_{ij}=x_{ij}$, and the design solves the following optimization problem:
@@
where $UE(s^{2})$ is defined in (3) and $c$ is the maximum number of compounds/well. As mentioned in the Introduction, we call the proposed row-constrained $UE(s^{2})$-optimal designs constrained row screening, or CRowS, designs.
@@ 2.1 Construction Algorithm
In order to construct approximately optimal row-constrained $UE(s^{2})$ designs, we use the coordinate exchange (CEx) heuristic (Meyer and Nachtsheim, 1995), along with some computational devices to make the procedure faster. Let $n_{1i}$ be the number of $+1$’s in design row $x_{i}$, so that $n_{1i}\leq c\;\forall i$. The algorithm is described in Algorithm 1. For the designs in this paper, we ran Algorithm 1 with 100 random starts and chose the design with the best objective function value.
@@ 2.2 Computational Improvements
The proof of this result is given in the Supplementary Materials. The lemma provides a simple test to determine if switching the signs of all $x_{ij}$ for fixed $i$ and $j\in J$ improves the criterion. The following result specializes it to the 1-coordinate and 2-coordinate exchanges in the proposed algorithm.
@@ 3 Properties of CRowS Designs
In this section, we study the designs we described in Section 2. In particular, we provide some initial bounds on the $UE(s^{2})$ criterion for the row-constrained designs, under the assumption that the row constraints are tight, which provides a design optimality check. Then, we empirically investigate several relevant design settings to better understand how the designs change as a function of the row constraint, providing insights into the practical effect of the constraints on the effectiveness of the row-constrained designs.
@@ 3.1 Some Bounds for CRowS Designs
In this section we adapt some ideas from Liu and Hickernell (2002). Recall the notation $Q(X):={\mathop{\mathrm{tr}}}(L^{{}^{\prime}}LL^{{}^{\prime}}L)$ and $L=[\mathbf{1},X]$ from §2.2. It is easily verified that $Q(X)=n^{2}+2\|X^{\prime}\mathbf{1}_{n}\|_{2}^{2}+{\mathop{\mathrm{tr}}}(X^{\prime}XX^{\prime}X)$.
@@ 3.2 Empirical Investigation of Row Constraint
An additional question, unanswered as of yet, is whether there is a threshold for $c$ below which we are guaranteed that a row-constrained $UE(s^{2})$-optimal design will have tight row constraints. In this section we investigate this question empirically. Reasoning intuitively, unconstrained $UE(s^{2})$-optimal designs try to make $s_{jl}$’s small, which means that the columns are pushed toward balance in order that the $s_{0j}$’s are small. In an unconstrained design, we therefore expect $+1$’s and $-1$’s should each occur approximately $nk/2$ times. Thus, for $c\geq k/2$, it is possible for the numbers of $+1$’s and $-1$’s to occur the same number of times, and we may observe slack in constraint (6) with the optimal $UE(s^{2})$ value similar to the unconstrained case. For $c<k/2$, however, we may see tight row constraints and, as $c$ gets smaller, an increasingly large difference between the row-constrained optimal $UE(s^{2})$ value and that of the unconstrained optimal design. That is, if $U^{*}$ is the $UE(s^{2})$ value for the unconstrained $UE(s^{2})$-optimal design and $U_{c}^{*}$ is the $UE(s^{2})$ value of the row-constrained $UE(s^{2})$-optimal design with $c$ $+1$’s per row, then $U^{*}\leq U_{c}^{*}$ if $c\leq k/2$, but as $c\rightarrow k$, $U_{c}^{*}\rightarrow U^{*}$.
@@
To explore the relationship between $c$ and $UE(s^{2})$, along with the effect of $c$ on the tightness of the constraints in (6), we first examine a set of approximately $UE(s^{2})$-optimal designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$, each constructed using $100$ algorithm tries of the algorithm described in Section 2.1. We consider several properties of the designs, each as a function of $c$: the $\text{UE}(s^{2})$ criterion value, along with the minimum, mean, and maximum slack in the row constraint, measured for each row as $(2c-k)-\sum_{j=1}^{k}x_{ij}$. We also plot the bounds from Section 3.1 for those designs without row constraint slack, and we provide some simulation results for various effect sizes. For the simulations, we generate $1{,}000$ datasets from model (1) with $\sigma^{2}=1$, $a=1$ nonzero element of $\bm{\beta}$ with value $\beta$, in line with application-specific sparsity levels of 1% or less. Since in these applications we know the desired effect directions, we consider $\beta\in\{0.375,0.5,0.75,1\}$, which translate to mean response differences (when the active compound is present versus absent) of $0.75$, $1$, $1.5$, and $2$, respectively. The datasets are analyzed using the Lasso as in (2). To be precise, we center $\mathbf{y}$ as $\mathbf{y}_{c}$ while also centering and scaling $X$ as $X_{cs}$ so that each column has mean 0 and the same length. The resulting Lasso estimates are denoted $\hat{\bm{\beta}}_{cs}$ with regularization parameter $\lambda_{cs}$. To reduce false positive rates, we perform a thresholding procedure in the following way. We assume known $\sigma^{2}$ (see discussion in Section 6 for a relaxation of this assumption) and, after obtaining $\hat{\bm{\beta}}$ for all $\lambda_{cs}$, we set any elements of $\hat{\bm{\beta}}_{cs}$ to 0 if they are less than $\sigma/8$, a value was chosen via testing to control FPR. For the resulting set of nonzero Lasso estimates, we obtain ordinary least squares estimates and choose the $\lambda_{cs}$ to minimize the Bayesian Information Criterion. The resulting set of nonzero estimates of $\bm{\beta}_{cs}$ constitute the set of factors deemed potential hits. The true positive rate (TPR) is the proportion of truly active factors that are correctly identified, and the false positive rate (FPR) is the proportion of truly inactive factors that are incorrectly deemed to be hits.
@@
Figure 1 provides a visualization of $\sqrt{UE(s^{2})}$ and the measures of row-constraint slack for all $c$ (left) and for a smaller set of $c$ (right). In Figure 2, we provide simulation results as a function of $c$. We observe that between $c=60$ and $c=70$ the constraint ceases to measurably impact the criterion value, the amount of row-constraint slack, and the TPR. We also see that while the bounds appear to be fairly tight (Figure 1, top left), closer inspection (Figure 1, top right) shows that there is a substantial gap as $c$ approaches the value beyond which there will be row-constraint slack. Thus, we conclude that for these designs any row constraint larger than $k/2=72$ will not have an appreciable effect on the results of an experiment using them, though the constraint includes slack beginning at some value $c$ slightly smaller than $k/2$. We have seen this same pattern for a number of other design sets as well (see Supplementary Material). The simulation (Figure 2) also shows that even for $c$ as small as 5 or 10, the designs reliably categorize active and inactive factors for sufficiently large effect sizes.
@@
Figure 1: Characteristics of CRowS designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$. Plots are a function of $c$, the number of $+1$’s allowed in each row of design. Top left: $UE(s^{2})$ value for each design (blue) along with the bounds from Section 3.1 (for all designs with no row-constraint slack). Bottom left: $(2c-k)-\sum_{j=1}^{k}x_{ij}$; the solid line is the average slack, while the lower and upper dashed lines are the minimum and maximum slack. Top right: The same as the top left plot, for $40\leq c\leq 84$. Bottom right: The same as the bottom left plot, for $40\leq c\leq 84$.
@@
Figure 2: Simulation results for CRowS designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$, for four effect sizes, assuming one active factor; the solid lines are true positive rates while the dashed lines are false positive rates.

### v2

@@ Abstract
High-throughput screening, in which large numbers of compounds are traditionally studied one-at-a-time in multiwell plates against specific targets, is widely used across many areas of the biological sciences including drug discovery. To improve the efficiency of these screens, we propose a new class of supersaturated designs that guide the construction of pools of compounds in each well. Because the size of the pools are typically limited by the particular application, the new designs accommodate this constraint and are part of a larger procedure that we call Constrained Row Screening or CRowS. We develop an efficient computational procedure to construct the CRowS designs, provide some initial lower bounds on the average squared off-diagonal values of their main-effects information matrix, and study the impact of the constraint on design quality. We also show via simulation that CRowS is statistically superior to the traditional one-compound-one-well approach as well as an existing pooling method, and demonstrate the use of the new methodology on a Verona Integron-encoded Metallo-$\beta$-lactamase 2 assay.
Keywords: experimental design; biological screening; the Lasso
@@ 1 Introduction
High-throughput screening (HTS) of chemical or biological compounds is used extensively across a range of critical scientific enterprises. This type of experiment is common in basic functional genomics studies and synthetic biology applications (Zeng et al., 2020; Sarnaik et al., 2020), but arguably its highest impact application is drug discovery (Blay et al., 2020; Soheilmoghaddam et al., 2021; Aldewachi et al., 2021). The literature suggests that nearly one-third of clinical drug candidates emerge from high-throughput screening campaigns (Brown and Boström, 2018), and these screens were a centerpiece in the National Cancer Institute’s Cancer Moonshot program (Wilson et al., 2020). For example, HTS is used in the search for inhibitors of enzymes that cause antibiotic resistance. The prevalence of such enzymes is of great concern because they exist in bacteria that cause nosocomial infections, including in secondary bacterial infections contributing to an estimated 50% of COVID-19 deaths (Mojica et al., 2022; Chedid et al., 2021). As liquid handling, miniaturization, and other screening technologies have improved, the number of potentially inhibitive compounds to explore has increased as well (Elkin et al., 2015; Volochnyuk et al., 2019), which motivates the need to improve the experimental and statistical aspects of these screens.
Classically, a single compound is applied to a single well in a 96-, 384-, or 1536-well plate (Malo et al., 2006; Kainkaryam and Woolf, 2009), an assay is performed with relevant normalization and bias-correction (Altman, 2005; Birmingham et al., 2009; Malo et al., 2010; Dragiev et al., 2011; Caraus et al., 2015), and hit compounds are identified with a z-score. As an alternative to one-compound-one-well (OCOW), researchers have suggested creating pools of compounds in each well which can improve HTS because compounds are observed in multiple wells rather than just one. Various versions of pooling have been used successfully (e.g., Devlin et al., 1996; Wilson-Lingardo et al., 1996; motlekar2008evaluation; Kainkaryam and Woolf, 2009; Elkin et al., 2015; silva2016output), and the most aggressive version, called non-adaptive pooling (Thierry-Mieg, 2006; Kainkaryam and Woolf, 2009), is said to accurately screen large numbers of compounds with relatively few pools. However, these methods are largely non-statistical in both their design and analysis.
In this work, we develop a novel class of row-constrained supersaturated designs that provides the pooling instructions to run a $k$-compound screen in $n$ wells, while accounting for a screener-specified constraint on the maximum number of compounds per well. The constraint may be due, for instance, to the size of the well, a required minimum compound concentration, or the maximum ionic strength of the solution beyond which the compounds may precipitate. We call these limitations row constraints, denoted by $c$, because they limit the number of $+1$’s in each design row. Designs are constructed using the unconditional $E(s^{2})$ criterion from the supersaturated design literature, and we analyze them using the Lasso. These designs are part of a larger procedure, which we call Constrained Row Screening, or CRowS, that includes both the design and its analysis via the Lasso.
We demonstrate that altogether CRowS is more flexible, cost-effective, and statistically efficient than traditional methods. We present a computationally efficient heuristic to construct the designs of any relevant $(n,k,c)$ combination, study the designs both theoretically and empirically as a function of $c$, provide evidence that our methodology improves upon both the OCOW approach as well as the pooling method of Kainkaryam and Woolf (2008), and show in a set of real experiments that it can be used to identify an inhibitor of a particular antibiotic-resistant enzyme. Thus, our contributions are a statistically principled and improved solution to an important biological problem and innovation in the construction of the relevant designs. (We note that after our work was completed we became aware of a recent paper by ji2023target that constructs pools by optimizing a “sensing matrix” which also uses a version of the Lasso to identify hits. They don’t explicitly constrain the size of the pools, nor do they engage with the statistical literature or study their method theoretically or via simulation. Still, it is encouraging that the biomedical community is considering statistical approaches to solve high-throughput screening problems.)
Our work emerges from research on supersaturated design (SSD), where $k$ factors are investigated using $n$ runs (with $n\leq k$). SSDs, which were first described over sixty years ago (Satterthwaite, 1959; Booth and Cox, 1962), have seen extensive development over the last thirty years (e.g., Lin, 1993; Wu, 1993; Nguyen, 1996; Cheng, 1997; Georgiou, 2014; Jones and Majumdar, 2014). In our setting, an SSD is a two-level, $n\times k$ design $X$, composed of $k$ compounds (factors; absence with level $-1$; presence with level $+1$), and $n$ wells (runs) with $k\geq n$ (along with a constraint on the design rows, described below and in Section 2). The SSD provides a set of instructions regarding which compounds are included in which well, so one may improve the design by constructing it using a statistically relevant criterion. SSDs are classically constructed using the $E(s^{2})$ criterion (Booth and Cox, 1962; Nguyen, 1996) to create balanced columns, as close to mutually orthogonal as possible. The row constraint of Section 2 typically necessitates a design with more $-1$’s than $+1$’s, so we use the unconditional $E(s^{2})$ criterion (Marley and Woods, 2010; Jones and Majumdar, 2014; Weese et al., 2015) which has the same orthogonality goal without the requirement of column balance. Mathematically, let $L=[\mathbf{1},X]$, and $S=L^{{}^{\prime}}L$ with elements $s_{ij}$ and $O_{S}=\sum_{i<j}s_{ij}^{2}$. Then the unconditional $E(s^{2})$ criterion is
$UE(s^{2})=\frac{O_{S}}{k(k+1)},$ (1)
the average of the squared off-diagonal elements of $S$. In many cases $UE(s^{2})$-optimal designs can be constructed directly (Jones and Majumdar, 2014), but for our row-constrained versions, we construct the designs algorithmically; see Section 2.1. After the design is constructed and executed, it is analyzed assuming the main-effects model
@@
where $\bm{\beta}=(\beta_{1},\dots,\beta_{k})^{T}$ and $\bm{\epsilon}\sim N(\mathbf{0},\sigma^{2}I)$. Because $k\geq n$, ordinary least squares cannot be used to estimate factor effects and regularization methods are preferred for analysis. The Gauss-Dantzig selector (candes_tao2007) is effective (Phoa et al., 2009; Marley and Woods, 2010; Weese et al., 2015), while the closely related Lasso (Tibshirani, 1996; Meinshausen et al., 2007; Lounici, 2008; James et al., 2009; Asif and Romberg, 2010) achieves similar performance (see Draguljić et al., 2014). Stallrich et al. (2023) has also recently developed an optimal screening design framework based on the Lasso. Thus, to analyze our proposed designs we use the Lasso, which selects estimates according to
@@
The remainder of the paper is structured as follows. In Section 2, we describe CRowS designs in more detail, along with the algorithm we use to construct them. Section 3 provides initial lower bounds on the $UE(s^{2})$ values for these CRowS designs, as well as an empirical study that sheds light on their behavior as a function of the severity of the row constraint. In Section 4 we use a set of simulations to show the effectiveness of our proposed procedure compared to alternative methods. In Section 5 we provide more detail on the metallo-$\beta$-lactamase system, and describe the results of several screens as a real-world proof-of-concept of CRowS. We provide a discussion and conclusion in Section 6.
@@ 2 Row-Constrained Supersaturated Designs
In this section we describe the new class of supersaturated designs that we propose, along with the heuristic optimization procedure developed to construct them. As explained in Section 1, $UE(s^{2})$-optimal designs minimize the average of the squared off-diagonal elements of information matrix $S$. In the present setting, we propose a generalization of $UE(s^{2})$-optimal designs that impose a constraint on the number of $+1$’s in each row. Thus, using notation from Section 1, we say that an $n\times k$ design $X$ is a CRowS design if $n\leq k$, $[X]_{ij}=x_{ij}$, and the design solves the following optimization problem:
@@
where $UE(s^{2})$ is defined in (1) and $c$ is the maximum number of compounds/well.
@@ 2.1 Construction Algorithm
In order to construct approximately optimal row-constrained $UE(s^{2})$ designs, we use the coordinate exchange (CEx) heuristic (Meyer and Nachtsheim, 1995), along with some computational devices to make the procedure faster. Let $n_{1i}$ be the number of $+1$’s in design row $x_{i}$, so that $n_{1i}\leq c\;\forall i$. The algorithm construction is described in Algorithm 1. For the designs in this paper, we ran Algorithm 1 with 100 random starts and chose the design with the best objective function value.
@@ 2.2 Computational Improvements
The proof of this result is given in Section 1 of Supplementary Material A. The proposition provides a simple test to determine if switching the signs of all $x_{ij}$ for fixed $i$ and $j\in J$ improves the criterion. The following result specializes it to the 1-coordinate and 2-coordinate exchanges in the proposed algorithm.
@@ 3 Properties of CRowS Designs
In this section, we study the designs described in Section 2. In particular, we provide some initial bounds on the $UE(s^{2})$ criterion for the row-constrained designs, under the assumption that the row constraints are tight, which provides a design optimality check. Then, we empirically investigate several relevant design settings to better understand how the designs change as a function of the row constraint, providing insights into the practical effect of the constraints on the effectiveness of the row-constrained designs.
@@ 3.1 Some Bounds for CRowS Designs
The result in this section uses ideas adapted from Liu and Hickernell (2002); it is proven in Section 2 of Supplementary Material A. Recall the notation $Q(X):={\mathop{\mathrm{tr}}}(L^{{}^{\prime}}LL^{{}^{\prime}}L)$ and $L=[\mathbf{1},X]$ from §2.2. It is easily verified that $Q(X)=n^{2}+2\|X^{\prime}\mathbf{1}_{n}\|_{2}^{2}+{\mathop{\mathrm{tr}}}(X^{\prime}XX^{\prime}X)$.
@@ 3.2 Empirical Investigation of Row Constraint
An additional unanswered question is whether there is a threshold for $c$ below which we are guaranteed that a row-constrained $UE(s^{2})$-optimal design will have tight row constraints. In this section we investigate this question empirically. Reasoning intuitively, unconstrained $UE(s^{2})$-optimal designs try to make $s_{jl}$’s small, which means that the columns are pushed toward balance in order that the $s_{0j}$’s are small. In an unconstrained design, we therefore expect $+1$’s and $-1$’s should each occur approximately $nk/2$ times. Thus, for $c\geq k/2$, it is possible for the numbers of $+1$’s and $-1$’s to occur the same number of times, and we may observe slack in constraint (6) with the optimal $UE(s^{2})$ value similar to the unconstrained case. For $c<k/2$, however, we may see tight row constraints and, as $c$ gets smaller, an increasingly large difference between the row-constrained optimal $UE(s^{2})$ value and that of the unconstrained optimal design. That is, if $U^{*}$ is the $UE(s^{2})$ value for the unconstrained $UE(s^{2})$-optimal design and $U_{c}^{*}$ is the $UE(s^{2})$ value of the row-constrained $UE(s^{2})$-optimal design, then $U^{*}\leq U_{c}^{*}$ if $c\leq k/2$, but as $c\rightarrow k$, $U_{c}^{*}\rightarrow U^{*}$.
@@
To explore the relationship between $c$ and $UE(s^{2})$, along with the effect of $c$ on the tightness of the constraints in (6), we first examine a set of approximately $UE(s^{2})$-optimal designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$, each constructed using $100$ algorithm tries of the algorithm described in Section 2.1. We consider several properties of the designs, each as a function of $c$: the $\text{UE}(s^{2})$ criterion value, along with the minimum, mean, and maximum slack in the row constraint, where slack is measured for each row $i$ as $(2c-k)-\sum_{j=1}^{k}x_{ij}$. We also plot the bounds from Section 3.1 for those designs without row constraint slack, and provide some simulation results. For the simulations, we follow the approach we describe in Section 4, assuming $a=1$ nonzero element of $\bm{\beta}$ with value $\beta$, in line with application-specific sparsity levels of 1% or less. Since we know the desired effect directions, we consider $\beta\in\{0.375,0.5,0.75,1\}$, which translate to differences in mean response (pools with vs. pool without active compound) of $0.75$, $1$, $1.5$, and $2$, respectively. The datasets are analyzed using the Lasso as in Section 4.1.3.
@@
Figure 1 provides a visualization of $\sqrt{UE(s^{2})}$ and the measures of row-constraint slack for all $c$ (left) and for a smaller set of $c$ (right). In Figure 2, we provide simulation results as a function of $c$. We observe that between $c=60$ and $c=70$ the constraint ceases to measurably impact the criterion value, the amount of row-constraint slack, and the TPR. We also see that while the bounds appear to be fairly tight (Figure 1, top left), closer inspection (Figure 1, top right) shows that there is a substantial gap as $c$ approaches the value beyond which there will be row-constraint slack. Thus, we conclude that for these designs any row constraint larger than $k/2=72$ will not have an appreciable effect on the results of an experiment using them, though the constraint includes slack beginning at some value $c$ slightly smaller than $k/2$. We have seen the same pattern for a number of other design sets as well (see Section 3 of Supplementary Material A). The simulation (Figure 2) also shows that even for $c$ as small as 5 or 10, the designs reliably categorize active and inactive factors for sufficiently large effect sizes.
@@
Figure 1: Characteristics of CRowS designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$. Plots are a function of $c$, the number of $+1$’s allowed in each row of design. Top left: $UE(s^{2})$ value for each design (solid line) along with the bounds from Section 3.1 (for all designs with no row-constraint slack) (dashed line). Bottom left: $(2c-k)-\sum_{j=1}^{k}x_{ij}$; the solid line is the average slack, while the lower and upper dashed lines are the minimum and maximum slack. Top right: The same as the top left plot, for $40\leq c\leq 84$. Bottom right: The same as the bottom left plot, for $40\leq c\leq 84$.
@@
Figure 2: Simulation results for CRowS designs with $n=96$, $k=144$ and $c\in\{2,3,\ldots,144\}$, for four effect sizes, assuming one active factor; the heavier lines above are true positive rates while the lighter lines below are false positive rates.